myininaya (myininaya):

I have the following information: \[n=p^2q\] \[k \ge 4\] \[0

OpenStudy (anonymous):

i can't even read this what class?

myininaya (myininaya):

i made a mistake \[\sum_{i=0}^{k}\left(\begin{matrix} k \\ i \end{matrix}\right) x^{k-i}(ypq)^i \equiv x^k+kypqx^{k-1}\] I'm trying to follow a proof but i don't know how they got from that one part to the other part

OpenStudy (anonymous):

yeah well i cannot read what it says. good luck

myininaya (myininaya):

http://www.cacr.math.uwaterloo.ca/hac/ its in chapter 11 on page 473 at the bottom

myininaya (myininaya):

;p;

myininaya (myininaya):

lol*

myininaya (myininaya):

i wonder if james or zarkon could do it

myininaya (myininaya):

oh its zarkon lol

OpenStudy (anonymous):

good luck!

OpenStudy (kinggeorge):

ask me next semester when I've taken cryptography, and I might be able to help you, but right now, I haven't got a clue.

OpenStudy (zarkon):

only the first two values of i will matter (i=0,1) for i>1 you have a \((pq)^2\) part which is \(p^2qq=nq\) which is zero nod n

myininaya (myininaya):

i don't see this (pq)^2 part

myininaya (myininaya):

george what kind of math have you had?

OpenStudy (zarkon):

\[{k\choose i}x^{k-1}(ypq)^i\] for \(i\ge2\) \[{k\choose i}x^{k-1}(ypq)^i={k\choose i}x^{k-1}(y^i)(pq)^{i-2}p^2qq\] \[=\left({k\choose i}x^{k-1}(y^i)(pq)^{i-2}q\right)\times n\]

myininaya (myininaya):

ok i see

myininaya (myininaya):

you are so smart to notice the first two values will only matter since n mod n is like totally 0 thanks zarkon

OpenStudy (kinggeorge):

I've only had a year and a half of college level math so far, so that means, calc 1-3, lin. algebra, abstract algebra, and analysis

OpenStudy (zarkon):

no problem :)

myininaya (myininaya):

is there any number theory in abstract algebra i can't remember

OpenStudy (kinggeorge):

not very much. I'm taking intro to number theory, and cryptography next semester.

myininaya (myininaya):

i like cryptography

myininaya (myininaya):

number theory is so cool and interesting to me although my skills suck right now

OpenStudy (kinggeorge):

It certainly sounds fun. I'm definitely enjoying my abstract algebra class right now. We had the barest hints of some number theory in discrete math so I'm not completely helpless.

myininaya (myininaya):

i like these math puzzles

myininaya (myininaya):

differential equations is not as fun to me but i do like the remedial kind lol

myininaya (myininaya):

zarkon do you think there is hope for my brain?

OpenStudy (kinggeorge):

From my experience, I don't like differential equations very much. Granted, I've never had a full course in it, but it just isn't my thing.

OpenStudy (zarkon):

there is!

myininaya (myininaya):

i feel like i'm trapped inside with an apes' brain

OpenStudy (zarkon):

A very smart ape ;)

myininaya (myininaya):

thanks i take that as a compliment

OpenStudy (zarkon):

If you want a fun class...Take a course in measure theoretic probability theory.

myininaya (myininaya):

he says that because he is a stat nerd

OpenStudy (kinggeorge):

I don't think my school has that class...

myininaya (myininaya):

my doesn't either

OpenStudy (zarkon):

or a class in stochastic processes

OpenStudy (zarkon):

sometimes it will fall under the title of applied probability

myininaya (myininaya):

i think they might have that

myininaya (myininaya):

i always had a 10 foot pole when it came to statistics i only had to put it down when the course in stat was required

OpenStudy (kinggeorge):

I'm personally more of an algebra person. Stats was interesting at first, but then it got boring.

myininaya (myininaya):

There are some cool things though in stat i do admit

OpenStudy (zarkon):

There is plenty of really nice mathematics in statistics

OpenStudy (kinggeorge):

It was the same with calculus for me. Interesting at first, and then it just stopped interesting me.

OpenStudy (zarkon):

A lot of multivariate calculus and matrix theory

myininaya (myininaya):

zarkon can i ask another question

myininaya (myininaya):

this is my last problem with this proof...

OpenStudy (zarkon):

ok

myininaya (myininaya):

omg never i got it lol

myininaya (myininaya):

i mean nvm

myininaya (myininaya):

i was typing it all out then i realized oh yeah lol

OpenStudy (zarkon):

ok :)

myininaya (myininaya):

ok i feel good about this proof thanks for all your help

OpenStudy (zarkon):

good...no problem